### v3

# Learning to Solve Integer Linear Programs with Davis-Yin Splitting

Abstract
In many applications, a combinatorial problem must be repeatedly solved with similar, but distinct parameters. Yet, the parameters $w$ are not directly observed; only contextual data $d$ that correlates with $w$ is available. It is tempting to use a neural network to predict $w$ given $d$. However, training such a model requires reconciling the discrete nature of combinatorial optimization with the gradient-based frameworks used to train neural networks. When the problem in question is an Integer Linear Program (ILP), one approach to overcome this training issue is to consider a continuous relaxation of the combinatorial problem. While existing methods utilizing this approach have shown to be highly effective on small problems, they do not always scale well to large problems. In this work, we draw on ideas from modern convex optimization to design a network and training scheme which scales effortlessly to problems with thousands of variables. Our experiments verify the computational advantage our proposed method enjoys on two representative problems, namely the shortest path problem and the knapsack problem.
1 Introduction
Many high-stakes decision problems in healthcare Zhong & Tang (2021), logistics and scheduling Sbihi & Eglese (2010); Kacem et al. (2021), and transportation Wang & Tang (2021) can be viewed as a two step process. In the first step, one gathers data about the situation at hand. This data is used to assign a value (or cost) to outcomes arising from each possible action. The second step is to select the action yielding maximum value (alternatively, lowest cost). Mathematically, this can be framed as an optimization problem with a data-dependent cost function:
$x^{\star}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{X}}f(x;d).$ (1)
In this work, we focus on the case where $\mathcal{X}\subset\mathbb{R}^{n}$ is a finite constraint set and $f(x;d)=w(d)^{\top}x$ is a linear function. This class of problems is quite rich, containing the shortest path, traveling salesperson, and sequence alignment problems, to name a few. Given $w(d)$, solving (1) may be straightforward (e.g. shortest path) or NP-hard (e.g. traveling salesperson problem Karp (1972)). However, our present interest is settings where the dependence of $w(d)$ on $d$ is unknown. In such settings, it is intuitive to learn a mapping $w_{\Theta}(d)\approx w(d)$ and then solve
${x}_{\Theta}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{X}}w_{\Theta}(d)^{\top}x$ (2)
in lieu of $w(d)$. The observed data $d$ is called the context. As an illustrative running example, consider the shortest path prediction problem shown in Figure 1, which is studied in Berthet et al. (2020) and Pogančić et al. (2019).
At first glance, it may appear natural to tune the weights $\Theta$ so as to minimize the difference between $w_{\Theta}(d)$ and $w(d)$. However, this is only possible if $w(d)$ is available at training time. Even if this approach is feasible, it may not be advisable. If $w_{\Theta}(d)$ is a near-perfect predictor of $w(d)$ we can expect $x_{\Theta}(d)\approx x^{\star}(d)$. However, if there are even small differences between $w_{\Theta}(d)$ and $w(d)$ this can manifest in wildly different solutions Bengio (1997); Wilder et al. (2019). Thus, we focus on methods where $\Theta$ is tuned by minimizing the discrepancy between $x_{\Theta}(d)$ and $x^{\star}(d)$.
The key obstacle in using gradient-based algorithms to tune $\Theta$ in such approaches is “differentiating through” the solution $x_{\Theta}(d)$ of (2) to obtain a gradient with which to update $\Theta$. Specifically, the combinatorial nature of $\mathcal{X}$ may cause the solution ${x}_{\Theta}(d)$ to remain unchanged for many small perturbations to $\Theta$; yet, for some perturbations ${x}_{\Theta}(d)$ may “jump” to a different point in $\mathcal{X}$. Hence, the gradient $\mbox{d}{x}_{\Theta}\big{/}\mbox{d}w_{\Theta}$ is always either zero or undefined. To compute an informative gradient, we follow Wilder et al. (2019) by relaxing (2) to a quadratic program over the convex hull of $\mathcal{X}$ with an added regularizer (see (8)).
Refer to caption Refer to caption Refer to caption
Figure 1: The shortest path prediction problem Pogančić et al. (2019). The goal is to find the shortest path (from top-left to bottom-right) through a randomly generated terrain map from the Warcraft II tileset Guyomarch . The contextual data $d$, shown in (a), is an image sub-divided into 8-by-8 squares, each representing a vertex in a 12-by-12 grid graph. The cost of traversing each square, i.e. $w(d)$, is shown in (b), with darker shading representing lower cost. The true shortest path is shown in (c).
Contribution
Drawing upon recent advances in convex optimization Ryu & Yin (2022) and implicit neural networks Fung et al. (2022); McKenzie et al. (2021), we propose a method designed specifically for “differentiating through” large-scale integer linear programs. Our approach is fast, easy to implement using our provided code, and, unlike several prior works (e.g. see Pogančić et al. (2019); Berthet et al. (2020)), trains completely on GPU. Numerical examples herein demonstrate our approach, run using only standard computing resources, easily scales to problems with tens of thousands of variables. Theoretically, we verify our approach computes an informative gradient via a refined analysis of Jacobian-free Backpropagation (JFB) Fung et al. (2022). In summary, we do the following.
[…]
    Building upon McKenzie et al. (2021), we use the three operator splitting technique of Davis & Yin (2017) to propose DYS-Net.
[…]
    We provide theoretical guarantees for differentiating through the fixed point of a certain non-expansive, but not contractive, operator.
[…]
    We numerically show DYS-Net easily handles combinatorial problems with tens of thousands of variables.
2 Preliminaries
LP Reformulation
We focus on optimization problems of the form (1) where $f(x;d)=w(d)^{\top}x$ and $\mathcal{X}$ is the integer or binary points of a polytope, which we may assume to be expressed in standard form Ziegler (2012):
$\mathcal{X}=\mathcal{C}\cap\mathbb{Z}^{n}\ \ \text{ or }\ \ \mathcal{X}=\mathcal{C}\cap\{0,1\}^{n},\quad\text{ where }\quad\mathcal{C}=\left\{x\in\mathbb{R}^{n}:\ Ax=b\text{ and }x\geq 0\right\}.$ (3)
In other words, (1) is an Integer Linear Program (ILP). Similar to works by Elmachtoub & Grigas (2022); Wilder et al. (2019); Mandi & Guns (2020) and others we replace the model (2) with its continuous relaxation, and redefine
${x}_{\Theta}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{C}}w_{\Theta}(d)^{\top}x$ (4)
as a step towards making the computation of $\mbox{d}{x}_{\Theta}\big{/}\mbox{d}w_{\Theta}$ feasible.
Losses and Training Data
We assume access to training data in the tuple form $(d,x^{\star}(d))$. We focus on the $\ell_{2}$ loss:
$\mathcal{L}(\Theta)\triangleq\mathbb{E}_{d\sim\mathcal{D}}\left[\ell_{2}(\Theta;d)\right],\\ \quad\text{ where }\quad\ell_{2}(\Theta;d)=\|x^{\star}(d)-{x}_{\Theta}(d)\|^{2},$ (5)
and $\mathcal{D}$ is the distribution of contextual data, although more sophisticated losses can be used with only superficial changes to our results. In principle we select the optimal weights by solving $\operatorname*{arg\,min}_{\Theta}\mathcal{L}(\Theta)$. In practice, the population risk is inaccessible, and so we minimize empirical risk instead Vapnik (1999):
$\operatorname*{arg\,min}_{\Theta}\frac{1}{N}\sum_{i=1}^{N}\ell_{2}\left(\Theta;d_{i}\right).$ (6)
Argmin Differentiation
For notational brevity, we temporarily suppress the dependence on $d$. The gradient of $\ell_{2}$ with respect to $\Theta$, evaluated on a single data point is
$\frac{\mathrm{d}}{\mathrm{d}\Theta}\left[\ell_{2}(\Theta)\right]=\frac{\mathrm{d}}{\mathrm{d}\Theta}\left[\|x_{\Theta}-x^{\star}\|^{2}\right]=(x_{\Theta}-x^{\star})^{\top}\frac{\partial x_{\Theta}}{\partial w_{\Theta}}\frac{\mathrm{d}w_{\Theta}}{\mathrm{d}\Theta}.$
As discussed in Section 1, ${x}^{\star}_{\Theta}$ is piecewise constant as a function of $w_{\Theta}$, and this remains true for the LP relaxation (4). Consequently, for all $w_{\Theta}$, either $\partial{x}_{\Theta}/\partial\Theta=0$ or $\partial{x}_{\Theta}/\partial\Theta$ is undefined—neither case yields an informative gradient. To remedy this, Wilder et al. (2019); Mandi & Guns (2020) propose adding a small amount of regularization to the objective function in (4) to make it strongly convex. This ensures ${x}_{\Theta}$ is a continuously differentiable function of $w_{\Theta}$. We add a small quadratic regularizer, modulated by $\gamma\geq 0$, to obtain the objective
$f_{\Theta}(x;\gamma,d)\triangleq w_{\Theta}(d)^{\top}x+\gamma\|x\|_{2}^{2}$ (7)
and henceforth replace (4) by
${x}_{\Theta}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{C}}f_{\Theta}(x;\gamma,d).$ (8)
During training, we aim to solve (8) and efficiently compute the derivative $\partial{x}_{\Theta}/\partial\Theta$.
3 Related Works
We highlight recent works combining optimization—particularly mixed integer programs—and deep learning.
Optimization Layers
Introduced in Amos & Kolter (2017), and studied further in Agrawal et al. (2019b; c); Bolte et al. (2021); Blondel et al. (2022), an optimization layer is a modular component of a deep learning framework where forward propagation consists of solving a parametrized optimization problem. Consequently, backpropagation entails differentiating the solution to this problem with respect to the parameters. Optimization layers are a promising technology as they are able to incorporate hard constraints into deep neural networks. Moreover, as their output is a (constrained) minimizer of the objective function, it is easier to interpret than the output of a generic layer.
Deep Equilibrium Models
Bai et al. (2019); El Ghaoui et al. (2021) propose the notion of deep equilibrium model (DEQ), also known as an implicit neural network. A DEQ is a neural network for which forward propagation consists of (approximately) computing a fixed point of a parametrized operator. We note that (8) may be reformulated as a fixed point problem,
Find $\text{Find }{x}_{\Theta}\text{ such that }{x}_{\Theta}=P_{\mathcal{C}}\left({x}_{\Theta}-\alpha\nabla_{x}f_{\Theta}({x}_{\Theta};d)\right),$ (9)
where $P_{\mathcal{C}}$ is the orthogonal projection [^1] For a set $\mathcal{A}\subseteq{\mathbb{R}}^{n}$, $P_{\mathcal{A}}(x)\triangleq\operatorname*{arg\,min}_{z\in\mathcal{A}}\|z-x\|$. onto $\mathcal{C}$. Thus, DEQ techniques may be applied to constrained optimization layers Chen et al. (2021); Blondel et al. (2022); McKenzie et al. (2021). However, the cost of computing $P_{\mathcal{C}}$ can be prohibitive, see the discussion in Section 4.
Differentiable Combinatorial Optimizers
We now narrow our focus to optimization layers for (integer) linear programs. Wilder et al. (2019) and Elmachtoub & Grigas (2022) are two of the earliest works applying deep learning techniques to data-driven LPs, and delineate two fundamentally different approaches. Specifically, Wilder et al. (2019) proposes replacing the LP with a continuous and strongly convex relaxation, as described in Section 2. This approach is extended to ILPs in Ferber et al. (2020), and to non-quadratic regularizers in Mandi & Guns (2020). On the other hand, Elmachtoub & Grigas (2022) avoid modifying the underlying optimization problem and instead propose a new loss function; dubbed the Smart Predict-then-Optimize (SPO) loss, to be used instead of the $\ell_{2}$ loss. We emphasize that the SPO loss requires access to the true cost vectors $w(d)$, a setting which we do not consider.
Several works define a continuously differentiable proxy for the solution to the unregularized LP (4), which we rewrite here as [^2] For sake of notational simplicity, the dependence on $d$ is implicit.
$x^{\star}(w)=\operatorname*{arg\,min}_{x\in\mathcal{C}}w^{\top}x,$ (10)
In Berthet et al. (2020), a stochastic perturbation is considered:
$x^{\star}_{\varepsilon}(w)=\mathbb{E}_{Z}\left[\operatorname*{arg\,min}_{x\in\mathcal{C}}\left(w+\varepsilon Z\right)^{\top}x\right],$ (11)
which is somewhat analogous to Nesterov-Spokoiny smoothing Nesterov & Spokoiny (2017) in zeroth-order optimization. Pogančić et al. (2019) define a piecewise-affine interpolant to $\ell(x^{\star}(w))$, where $\ell$ is a suitable loss function.
We note a bifurcation in the literature; Wilder et al. (2019); Ferber et al. (2020); Elmachtoub & Grigas (2022) assume access to training data of the form $(d,w(d))$, whereas Pogančić et al. (2019); Berthet et al. (2020); Sahoo et al. (2022) use training data of the form $(d,x^{\star}(d))$. Our focus is on the latter setting, and we refer the reader to Kotary et al. (2021) for a recent survey of this area.
Learning-to-Optimize (L2O)
Our work can be situated within the broader learning-to-optimize (L2O) framework Chen et al. (2022); Li & Malik (2017); Chen et al. (2018); Liu et al. (2023), where an optimization algorithm is learned and its outputs are used to perform inferences. However, traditional L2O methods use a fixed number of layers (i.e. unrolled algorithms). Our approach attempts to differentiate through the solution $x^{\star}_{\Theta}$ and is therefore most closely aligned with works that employ implicit networks within the L2O framework Amos & Kolter (2017); Heaton & Fung (2023); McKenzie et al. (2021); Gilton et al. (2021); Liu et al. (2022). We highlight that, unlike many L20 works, our goal is not to learn a faster optimization method for fully specified problems. Rather, we seek to solve partially specified problems given contextual data by combining learning and optimization techniques.
Computing the derivative of a minimizer with respect to parameters
In all of the aforementioned works, the same fundamental problem is encountered: $\partial{x}_{\Theta}/\partial\Theta$ must be computed where $x_{\Theta}$ is the solution of a (constrained) optimization problem with objective function parametrized by $\Theta$. The most common approach to doing so, proposed in Amos & Kolter (2017) and used in Ruthotto et al. (2018); Wilder et al. (2019); Mandi & Guns (2020); Ferber et al. (2020), starts with the KKT conditions for constrained optimality:
$\frac{\partial f_{\Theta}}{\partial x}({x}_{\Theta})+A^{\top}\hat{\lambda}+\hat{\nu}=0,\ A{x}-b=0,\ D(\hat{\nu}){x}_{\Theta}=0,$
where $\hat{\lambda}$ and $\hat{\nu}\geq 0$ are Lagrange multipliers associated to the optimal solution ${x}_{\Theta}$ Bertsekas (1997) and $D(\hat{\nu})$ is a matrix with $\hat{\nu}$ along its diagonal. Differentiating these equations with respect to $\Theta$ and rearranging yields
$\begin{bmatrix}\frac{\partial^{2}f_{\Theta}}{\partial x^{2}}&A&I\\[4.0pt] A^{\top}&0&0\\[4.0pt] D(\hat{\nu})&0&D({x}_{\Theta})\end{bmatrix}\begin{bmatrix}\frac{d{x}_{\Theta}}{d\Theta}\\[4.0pt] \frac{d\hat{\lambda}}{d\Theta}\\[4.0pt] \frac{d\hat{\nu}}{d\Theta}\end{bmatrix}=\begin{bmatrix}\frac{\partial^{2}f_{\Theta}}{\partial x\partial\Theta}\\[4.0pt] 0\\[4.0pt] 0\end{bmatrix}.$ (12)
The matrix and right hand side vector in (12) are computable, thus enabling one to solve for $\frac{d{x}_{\Theta}}{d\Theta}$ (as well as $\frac{d\hat{\lambda}}{d\Theta}$ and $\frac{d\hat{\nu}}{d\Theta}$). The computational bottleneck in this approach is computing the Lagrange multipliers at optimality—i.e. $\hat{\lambda}$ and $\hat{\nu}$—in addition to ${x}_{\Theta}$. If $x\in{\mathbb{R}}^{n}$ and $A\in\mathbb{R}^{m\times n}$ this can be done with a primal-dual interior point method at a cost of $\mathcal{O}\left(\max\{n,m\}^{3}\right)$ Amos & Kolter (2017).
Using the stochastic proxy (11), Berthet et al. (2020) derive a formula for $dx^{\star}_{\varepsilon}/dw$ which is also an expectation, and hence can be efficiently approximated using Monte Carlo methods. Pogančić et al. (2019) show that the gradients of their interpolant are strikingly easy to compute, requiring just one additional solve of (10) with perturbed cost $w^{\prime}$. This approach is extended by Sahoo et al. (2022) which proposes to avoid computing $dx^{\star}_{\varepsilon}/dw$ entirely, replacing it with the negative identity matrix. This is similar in spirit to our use of Jacobian-free Backpropagation (see Section 4).
4 DYS-Net
We now introduce our proposed model, DYS-net. We use this term to refer to the model and the training procedure. Fixing an architecture for $w_{\Theta}$, and an input $d$, DYS-net computes an approximation to $x_{\Theta}(d)$ in a way that is easy to backpropagate through:
[…]
The Forward Pass
As we wish to compute ${x}_{\Theta}$ and $\partial{x}_{\Theta}/\partial\Theta$ for high dimensional settings (i.e. large $n$), we eschew second-order methods (e.g. Newton’s method) in favor of first-order methods such as projected gradient descent (PGD). With PGD, a sequence $\{x^{k}\}$ of estimates of $x_{\Theta}$ are computed so that $x_{\Theta}=\lim_{k\rightarrow\infty}x^{k}$ where
$x^{k+1}=P_{\mathcal{C}}\left(x^{k}-\alpha\nabla_{x}f(x^{k};\gamma,d)\right)\quad\text{for all}\ k\in{\mathbb{N}}.$
This approach works for simple sets $\mathcal{C}$ for which there exists an explicit form of $P_{\mathcal{C}}$, e.g. when $\mathcal{C}$ is the probability simplex Duchi et al. (2008); Condat (2016); Li et al. (2023). However, for general polytopes $\mathcal{C}$ no such form exists, thereby requiring a second iterative procedure, run at each iteration $k$, to compute $P_{\mathcal{C}}(x^{k})$. To avoid this computational expense, we draw on recent advances in the convex optimization literature [^3] See Pedregosa & Gidel (2018); Yurtsever et al. (2021) for similar applications of DYS, particularly Davis-Yin splitting or DYS Davis & Yin (2017). Specifically, we adapt the architecture incorporating DYS proposed in McKenzie et al. (2021). To this end, we rewrite $\mathcal{C}$ as
$\mathcal{C}=\{x:\ Ax=b\text{ and }x\geq 0\}=\underbrace{\{x:\ Ax=b\}}_{\triangleq\,\mathcal{C}_{1}}\cap\underbrace{\{x:\ x\geq 0\}}_{\triangleq\,\mathcal{C}_{2}}=\mathcal{C}_{1}\cap\mathcal{C}_{2}.$ (14)
While $P_{\mathcal{C}}$ is hard to compute, both $P_{\mathcal{C}_{1}}$ and $P_{\mathcal{C}_{2}}$ can be computed cheaply via explicit formulas, once a singular value decomposition (SVD) is computed for $A$. We verify this via the following lemma (included for completeness, as the two results are already known).
Lemma 1.
If $\mathcal{C}_{1},\mathcal{C}_{2}$ are as in (14) and $A$ is full-rank then:
  1. 1.
    $P_{\mathcal{C}_{1}}(z)=z-A^{\dagger}(Az-b)$, where $A^{\dagger}=V\Sigma^{-1}U^{\top}$ and $U\Sigma V^{\top}$ is the compact SVD of $A$.
  2. 2.
    $P_{\mathcal{C}_{2}}(z)=\operatorname{ReLU}(z)\triangleq\max\{0,z\}$.
The following theorem allows one to approximate $x_{\Theta}$ using only $P_{\mathcal{C}_{1}}$ and $P_{\mathcal{C}_{2}}$, not $P_{\mathcal{C}}$.
Theorem 2.
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be as in (14), and suppose $f_{\Theta}(x;\gamma,d)=w_{\Theta}(d)^{\top}x+\frac{\gamma}{2}\|x\|_{2}^{2}$ for any neural network $w_{\Theta}(d)$. For any $\alpha\in(0,2/\gamma)$ define the sequence $\{z^{k}\}$ by:
$z^{k+1}=T_{\Theta}(z^{k})\quad\text{ for all }k\in\mathbb{N}$ (15)
where
$T_{\Theta}(z)\!\!\triangleq z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\!\left(\left(2-\alpha\gamma\right)P_{\mathcal{C}_{2}}(z)\!-\!z\!-\!\alpha w_{\Theta}(d)\right).$ (16)
If $x^{k}\triangleq P_{\mathcal{C}_{2}}(z^{k})$ then $\|x^{k+1}-x^{k}\|^{2}_{2}=O(1/k)$.
A full proof is included in Appendix A; here we provide a proof sketch. Substituting the gradient
$\nabla_{z}f_{\Theta}(z;\gamma,d)=w_{\Theta}(d)+\gamma z$
for $F_{\Theta}(\cdot,d)$ into the formula for $T_{\Theta}(\cdot)$ given in (McKenzie et al., 2021, Theorem 3.2) and rearranging yields (16). Verifying the assumptions of (McKenzie et al., 2021, Theorem 3.3) are satisfied is straightforward. For example, $\nabla_{z}f_{\Theta}(z;\gamma,d)$ is $1/\gamma$-cocoercive by the Baillon-Haddad theorem Baillon & Haddad (1977); Bauschke & Combettes (2009).
The forward pass of DYS-net consists of iterating (15) until a suitable stopping criterion is met, which for simplicity we assume to be a maximum number of iterations.
The Backward Pass
From Theorem 2 we deduce the following fixed point condition:
${x}_{\Theta}=P_{\mathcal{C}_{2}}({z}_{\Theta}),\quad\text{ for some }\ {z}_{\Theta}\in\{z:z=T_{\Theta}({z})\}.$ (17)
As discussed in McKenzie et al. (2021), instead of backpropagating through every iterate of the forward pass, we may derive a formula for $dx_{\Theta}/d\Theta$ by appealing to the implicit function theorem and differentiating both sides of (17):
$\frac{\mathrm{d}{z}_{\Theta}}{\mathrm{d}\Theta}=\frac{\partial T_{\Theta}}{\partial\Theta}+\frac{\partial T_{\Theta}}{\partial z}\frac{\mathrm{d}{z}_{\Theta}}{\mathrm{d}\Theta}\quad\implies\quad\mathcal{J}_{\Theta}(z_{\Theta})\ \frac{\mathrm{d}{z}_{\Theta}}{\mathrm{d}\Theta}=\frac{\partial T_{\Theta}}{\partial\Theta}\text{ where }\mathcal{J}_{\Theta}(z)=I-\frac{\partial T_{\Theta}}{\partial z}.$ (18)
We notice two immediate problems not addressed by McKenzie et al. (2021): (i) $T_{\Theta}$ is not everywhere differentiable with respect to $z$, as $P_{\mathcal{C}_{2}}$ is not; (ii) if $T_{\Theta}$ were a contraction (i.e. Lipschitz in $z$ with constant less than $1$), then $\mathcal{J}_{\Theta}$ would be invertible. However, this is not necessarily the case. Thus, it is not clear a priori that (18) can be solved for $\mbox{d}{{z}_{\Theta}}/\mbox{d}{\Theta}$. Our key result (Theorem 6 below) is to provide reasonable conditions under which $\mathcal{J}_{\Theta}(z_{\Theta})$ is invertible.
Assuming these issues can be resolved, one may compute the gradient of the loss using the chain rule:
$\frac{\mathrm{d}\ell}{\mathrm{d}\Theta}=\frac{\mathrm{d}\ell}{\mathrm{d}x}\frac{\mathrm{d}{x}_{\Theta}}{\mathrm{d}\Theta}=\frac{\mathrm{d}\ell}{\mathrm{d}x}\left(\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\frac{\mathrm{d}{z}_{\Theta}}{\mathrm{d}\Theta}\right)=\frac{\mathrm{d}\ell}{\mathrm{d}x}\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\mathcal{J}_{\Theta}^{-1}\frac{\partial T_{\Theta}}{\partial\Theta}$
This approach requires solving a linear system with $\mathcal{J}_{\Theta}$ which becomes particularly expensive when $n$ is large. Instead, we use Jacobian-free Backpropagation (JFB) Fung et al. (2022), also known as one-step differentiation Bolte et al. (2024), in which the Jacobian $\mathcal{J}_{\Theta}$ is replaced with the identity matrix. This leads to an approximation of the true gradient $\mbox{d}{\ell}/\mbox{d}{\Theta}$ using
$p_{\Theta}(x_{\Theta})=\left[\frac{\partial\ell}{\partial x}\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\frac{\partial T_{\Theta}}{\partial\Theta}\right]_{(x,z)=(x_{\Theta},z_{\Theta})}.$ (19)
We show (19) is a valid descent direction by resolving the two problems highlighted above. We begin by rigorously deriving a formula for $\partial{T_{\Theta}}/\partial{z}$. Recall the following generalization of the Jacobian to non-smooth operators due to Clarke (1983).
Definition 3.
For any locally Lipshitz $F:\mathbb{R}^{n}\to\mathbb{R}^{n}$ let $D_{F}$ denote the set upon which $F$ is differentiable. The Clarke Jacobian of $F$ is the set-valued function defined as
$\partial F(\bar{z})=\left\{\begin{array}[]{cc}\left.\frac{\mathrm{d}F}{\mathrm{d}z}\right|_{z=\bar{z}}&\text{ if }\bar{z}\in D_{F}\\ \operatorname{Con}\left\{\lim_{z^{\prime}\to\bar{z}:z^{\prime}\in D_{F}}\left.\frac{\mathrm{d}F}{\mathrm{d}z}\right|_{z=z^{\prime}}\right\}&\text{ if }\bar{z}\notin D_{F}\end{array}\right.$
Where $\operatorname{Con}\left\{\right\}$ denotes the convex hull of a set.
The Clarke Jacobian of $P_{\mathcal{C}_{2}}$ is easily computable, see Lemma 9. Define the (set-valued) functions
$c(\alpha)\triangleq\partial\max(0,\alpha)=\begin{cases}\begin{array}[]{cl}1&\text{ if }\alpha>0\\ 0&\text{ if }\alpha<0\\[0.1pt] &\text{ if }\alpha=0\end{array}\end{cases}\quad\text{ and }\quad\tilde{c}(\alpha)=\begin{cases}1&\text{ if }\alpha>0\\ 0&\text{ if }\alpha\leq 0\end{cases}$
Then
$\partial P_{\mathcal{C}_{2}}(\bar{z})=\left[\frac{\mathrm{d}}{\mathrm{d}z}\operatorname{ReLU}(z)\right]_{z=\overline{z}}=\mathrm{diag}(c(\overline{z})),$ (20)
where $c$ is applied element-wise. We shall now provide a consistent rule for selecting, at any $\bar{z}$, an element of $\partial P_{\mathcal{C}_{2}}(\bar{z})$. If $z_{i}\neq 0$ for all $i$ then $\partial P_{\mathcal{C}_{2}}$ is a singleton. If one or more $z_{i}=0$ then $\partial P_{\mathcal{C}_{2}}$ is multi-valued, so we choose the element of $\partial P_{\mathcal{C}_{2}}$ with 00 in the $(i,i)$ position for every $z_{i}=0$. We write $dP_{\mathcal{C}_{2}}/dz$ for this chosen element, and note that
$\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z=\bar{z}}=\mathrm{diag}(\tilde{c}(\overline{z}))\in\partial P_{\mathcal{C}_{2}}(\bar{z})$
This aligns with the default rule for assigning a sub-gradient to $\operatorname{ReLU}$ used in the popular machine learning libraries TensorFlowAbadi et al. (2016), PyTorch Paszke et al. (2019) and JAX Bradbury et al. (2018), and has been observed to yield networks which are more stable to train than other choices Bertoin et al. (2021).
Given the above convention, we can compute $\partial{T_{\Theta}}/\partial{z}$, interpreted as an element of the Clarke Jacobian of $T_{\Theta}$ chosen according to a consistent rule. Surprisingly, $\partial{T_{\Theta}}/\partial{z}$ may be expressed using only orthogonal projections to hyperplanes. Throughout, we let $e_{i}\in\mathbb{R}^{n}$ be the one-hot vector with $1$ in the $i$-th position and zeros elsewhere, and $a_{i}^{\top}$ be the $i$-th row of $A$. For any subspace $\mathcal{H}$ we denote the orthogonal subspace as $\mathcal{H}^{\perp}$. The following theorem, the proof of which is given in Appendix A, characterizes $\partial{T_{\Theta}}/\partial{z}$.
Theorem 4.
Suppose $A$ is full-rank and $\mathcal{H}_{1}\triangleq\operatorname{Null}(A)$, $\mathcal{H}_{2,z}\triangleq\operatorname{Span}\left(e_{i}:z_{i}>0\right)$. Then for all $\hat{z}\in{\mathbb{R}}^{n}$
$\left.\frac{\partial T_{\Theta}}{\partial z}\right|_{z=\bar{z}}=P_{\mathcal{H}_{1}^{\bot}}P_{\mathcal{H}_{2,\bar{z}}^{\bot}}+(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\bar{z}}}.$ (21)
Algorithm 1 DYS-Net
1:  Inputs:$A$ and $b$ defining $\mathcal{C}$, hyperparameters $\alpha,\gamma,K$
2:  Initialize: Attach neural network $w_{\Theta}(\cdot)$. Compute SVD of $A$ for $P_{\mathcal{C}_{1}}$ formula
3:  Forward Pass: Initialize $z^{0}$. Given $w_{\Theta}(d)$ compute $z^{1},\ldots,z^{K}$ using (15). Return $x^{K}\triangleq P_{\mathcal{C}_{1}}(z^{K})\approx{x}_{\Theta}$.
4:  Backward Pass: Compute $p_{\Theta}(x^{K})\approx p_{\Theta}({x}_{\Theta})$ using (19) and return.
To show JFB is applicable, it suffices to verify $\|\partial{T_{\Theta}}/\partial{z}\|<1$ when evaluated at ${z}_{\Theta}$. Theorem 4 enables us to show this inequality holds when ${x}_{\Theta}$ satisfies a commonly-used regularity condition, which we formalize as follows.
Definition 5 (LICQ condition, specialized to our case).
Let ${x}_{\Theta}$ denote the solution to (8). Let $\mathcal{A}({x}_{\Theta})\subseteq\{1,\ldots,n\}$ denote the set of active positivity constraints:
$\mathcal{A}({x}_{\Theta})\triangleq\{i:\ [{x}_{\Theta}]_{i}=0\}.$ (22)
The point ${x}_{\Theta}$ satisfies the Linear Independence Constraint Qualification (LICQ) condition if the vectors
$\{a_{1},\ldots,a_{m}\}\cup\{e_{i}:i\in\mathcal{A}({x}_{\Theta})\}$ (23)
are linearly independent.
Theorem 6.
If the LICQ condition holds at ${x}_{\Theta}$, $A$ is full-rank and $\alpha\in(0,2/\gamma)$, then $\|\partial T_{\Theta}/\partial z\|_{z={z}_{\Theta}}<1$.
The significance of Theorem 6, which is proved in Appendix A, is outlined by the following corollary, stating that using $p_{\Theta}$ instead of the true gradient $d\ell/d\Theta$ is still guaranteed to decrease $\ell(\Theta)$, at least for small enough step-size.
Corollary 7.
If $T_{\Theta}$ is continuously differentiable with respect to $\Theta$ at $z_{\Theta}$, the assumptions in Theorem 6 hold and $(\partial T_{\Theta}/\partial\Theta)^{\top}(\partial T_{\Theta}/\partial\Theta)$ has condition number sufficiently small, then $p_{\Theta}(x_{\Theta})$ is a descent direction for $\ell$ with respect to $\Theta$.
Theorem 6 also proves that $\mathcal{J}_{\Theta}(z_{\Theta})$ in (18) is invertible, thus also justifying the use of Jacobian-based backprop as well as numerous other gradient approximation techniques Liao et al. (2018); Geng et al. (2021).
Finally, we note that in practice $p_{\Theta}(x^{K})$ is used instead of $p_{\Theta}(x_{\Theta})$. (Fung et al., 2022, Corollary 0.1) guarantees that $p_{\Theta}(x^{K})$ will still be a descent direction for $K$ sufficiently large, see also (Bolte et al., 2024, Cor. 1).
Implementation
DYS-net is implemented as an abstract PyTorch Paszke et al. (2019) layer in the provided code. To instantiate it, a user only need specify $A$ and $b$ (see (3)) and choose an architecture for $w_{\Theta}$. At test time, it can be beneficial to solve the underlying combinatorial problem (2) exactly using a specialized algorithm, e.g. commercial integer programming software for the knapsack prediction problem. This can easily be done using DYS-net by instantiating the test-time-forward abstract method.
5 Numerical Experiments
5.1 Experiments using PyEPO
First, we verify the efficiency of DYS-net by performing benchmarking experiments within the PyEPO Tang & Khalil (2022) framework. We compare DYS-net against three other methods: Perturbed Optimization Berthet et al. (2020)PertOpt-net; an approach using Blackbox Backpropagation Vlastelica et al. (2019)BB-net; and an approach using cvxpylayers Agrawal et al. (2019a)CVX-net.
Models
We use the PyEPO implementations of Perturbed Optimization, respectively Blackbox Backpropagation, as the core of PertOpt-net, respectively BB-net. We use the same architecture for $w_{\Theta}(d)$ for all models [^4] The architecture we use for shortest path prediction problems differs from that used for knapsack prediction problems. See the appendix for details. and use an appropriate combinatorial solver at test time, via the PyEPO interface to Gurobi. Thus, the methods differ only in training procedure. The precise architectures used for $w_{\Theta}(d)$ for each problem are described in Appendix C, and are also viewable in the accompanying code repository [^5] See code at https://github.com/mines-opt-ml/fpo-dys .Interestingly, we observed that adding drop-out during training to the output layer proved beneficial for the knapsack prediction problem. Without this, we find that $w_{\Theta}$ tends to output a sparse approximation to $w$ supported on a feasible set of items, and so does not generalize well.
Training
We train for a maximum of 100 epochs or 30 minutes, whichever comes first. We use a validation set for model selection as we observe that, for all methods, the best loss is seldom achieved at the final iteration. Exact hyperparameters are discussed in the appendix. For each problem size, we perform three repeats with different randomly generated data. Results are displayed in Figure 2.
Evaluation
Given test data $\{(d_{i},w(d_{i}),x^{\star}(d_{i}))\}_{i=1}^{N}$ and a fully trained model $w_{\Theta}(d)$ we define the regret as:
$r(d)=w(d_{i})^{\top}\left(x_{\Theta}(d_{i})-x^{\star}(d_{i})\right).$ (24)
Note that regret is non-negative, and is zero precisely when $x_{\Theta}(d_{i})$ is a solution of equal quality to $x^{\star}(d_{i})$. Following Tang & Khalil (2022), we evaluate the quality of $w_{\Theta}(d)$ using normalized regret, defined as
$\tilde{r}=\frac{\sum_{i=1}^{N}r(d_{i})}{\sum_{i=1}^{N}\left[w(d_{i})^{\top}x^{\star}(d_{i})\right]}.$ (25)
5.1.1 Shortest Path Prediction
The shortest path between two vertices in a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ can be found by:
$x^{\star}=\operatorname*{arg\,min}_{x\in\mathcal{X}}w(d)^{\top}x;\ \mathcal{X}=\{x\in\{0,1\}^{|\mathcal{E}|}:Ex=b\},$ (26)
where $E$ is the vertex-edge adjacency matrix, $b$ encodes the initial and terminal vertices, and $w(d)\in\mathbb{R}^{|\mathcal{E}|}$ is a vector encoding ($d$-dependent) edge lengths; see Appendix for further details. Here, $x^{\star}$ will encode the edges included in the optimal path. In this experiment we focus on the case where $\mathcal{G}$ is the $k\times k$ grid graph. We use the PyEPO Tang & Khalil (2022) benchmarking software to generate training datasets of the form $\{(d_{i},x^{\star}_{i}\approx x^{\star}(d_{i}))\}_{i=1}^{N}$ where $N=1000$ for each $k\in\{5,10,15,20,25,30\}$. The $d_{i}$ are sampled from the five-dimensional multivariate Gaussian distribution with mean 00 and identity covariance. Note that the number of variables in (26) scales like $k^{2}$, not $k$.
5.1.2 Knapsack Prediction
In the (0–1, single) knapsack prediction problem, we are presented with a container (i.e. a knapsack) of size $c$ and $I$ items, of sizes $s_{1},\ldots,s_{I}$ and values $w_{1}(d),\ldots,w_{I}(d)$. The goal is to select the subset of maximum value that fits in the container, i.e. to solve:
$x^{\star}=\operatorname*{arg\,max}_{x\in\mathcal{X}}w(d)^{\top}x;\ \mathcal{X}=\{x\in\{0,1\}^{I}:\ s^{\top}x\leq c\}$
Here, $x^{\star}$ encodes the selected items. In the (0–1) $k$-knapsack prediction problem we imagine the container having various notions of “size” (i.e. length, volume, weight limit) and hence a $k$-tuple of capacities $c\in\mathbb{R}^{k}$. Correspondingly, the items each have a $k$-tuple of sizes $s_{1},\ldots,s_{I}\in\mathbb{R}^{k}$. We aim to select a subset of maximum value, amongst all subsets satisfying the $k$ capacity constraints:
$\begin{split}&x^{\star}=\operatorname*{arg\,max}_{x\in\mathcal{X}}w(d)^{\top}x\\ &\mathcal{X}=\{x\in\{0,1\}^{I}:\ Sx\leq c\}\\ &S=\begin{bmatrix}s_{1}&\cdots&s_{k}\end{bmatrix}\in\mathbb{R}^{k\times I}\end{split}$ (27)
In Appendix B we discuss how to transform $\mathcal{X}$ into the canonical form discussed in Section 2. We again use PyEPO to generate datasets of the form $\{(d_{i},x^{\star}_{i}\approx x^{\star}(d_{i}))\}_{i=1}^{N}$ where $N=1000$ for $k=2$ and $I$ varying from $50$ to $750$ inclusive in increments of $50$. The $d_{i}$ from the five-dimensional multivariate Gaussian distribution with mean 00 and identity covariance.
(a) Normalized Regret
(b) Train Time (in seconds)
(c) Normalized Regret
(d) Train Time (in seconds)
Figure 2: Results for the shortest path and knapsack prediction problems. Figures (a) and (b) show normalized regret and train time for the shortest path prediction problem, while Figures (c) and (d) show normalized regret and train time for the knapsack prediction problem. Note that the train time in figures (b) and (d) is the time till the model achieving best normalized regret on the validation set is reached.
5.1.3 Results
As is clear from Figure 2 DYS-net trains the fastest among the four methods, and achieves the best or near-best normalized regret in all cases.
5.2 Large-scale shortest path prediction
Our shortest path prediction experiment described in Section 5.1.1 is bottlenecked by the fact that the “academic” license of Gurobi does not allow for shortest path prediction problems on grids larger than $30\times 30$. To further explore the limits of DYS-net, we redo this experiment using a custom pyTorch implementation of Dijkstra’s algorithm666adapted from Tensorflow code available at https://github.com/google-research/google-research/blob/master/perturbations/experiments/shortest˙path.py as our base solver for (26).
Data Generation
We generate datasets $\mathcal{D}=\{(d_{i},x^{\star}(d_{i})\}_{i=1}^{1,000}$ for $k$-by-$k$ grids where $k\in\{5,10,20,30,50,100\}$ and the $d_{i}$ are sampled uniformly at random from $[0,1]^{5}$, the true edge weights are computed as $w(d)=Wd$ for fixed $W\in\mathbb{R}^{|\mathcal{E}|\times 5}$, and $x^{\star}(d)$ is computed given $w(d)$ using the aforementioned pyTorch implementation of Dijkstra’s algorithm. Further details are presented in Appendix C.
Figure 3: Accuracy (in percentage) of predicted paths on 5-by-5 grid during training.
grid size number of variables network size
5-by-5 40 500
10-by-10 180 2040
20-by-20 760 8420
30-by-30 1740 19200
50-by-50 4900 53960
100-by-100 19800 217860
Table 1: Number of variables (i.e. number of edges) per grid size for the shortest path prediction problem of Section 5.2. Third column: number of parameters in $w_{\Theta}(d)$ for DYS-Net, CVX-net and PertOpt-net. For BB-Net, we found a latent dimension that is 20-times larger than the aforementioned three to be more effective.
(a) Test MSE Loss
(b) Training Time (min)
(c) Regret Values
Figure 4: Results for the shortest path prediction problem. a) Test MSE loss (left), b) training time in minutes (middle), and c) regret values (right) vs. gridsize for DYS-Net (proposed) and approaches using cvxpylayers Agrawal et al. (2019b) labeled CVX; Perturbed Optimization Berthet et al. (2020) labeled PertOpt; and Blackbox Backpropagation Vlastelica et al. (2019), labeled BB. Note CVX is unable to load or run problems with gridsize over 30. Dimensions of the variables can be found in Table 1.
Models and Training
We test the same four approaches as in Sections 5.1.1 and 5.1.2, but unlike in these sections we do not use the PyEPO implementations of PertOpt-net or BB-net. We cannot, as the PyEPO implementations call Gurobi to solve (26) which, as mentioned above, cannot handle grids larger than $30\times 30$. Instead, we use custom implementations based on the code accompanying the works introducing PertOpt Berthet et al. (2020) and BB Vlastelica et al. (2019). We use the same architecture for $w_{\Theta}(d)$ for DYS-net, PertOpt-net, and Cvx-net; a two layer fully connected neural network with leaky ReLU activation functions. For DYS-net and Cvx-net we do not modify the forward pass at test time. For BB-net we use a larger network by making the latent dimension 20-times larger than that of the first three as we found this more effective. Network sizes can be seen in Table 1.
We tuned the hyperparameters for each approach to the best of our ability on the smallest problem (5-by-5 grid) and then used these hyperparameter values for all other graph sizes. See Figure 3 for the results of this training run. We train all approaches for 100 epochs total on each problem using the $\ell_{2}$ loss (5).
Results
The results are displayed in Figure 4. While CVX-net and PertOpt-net achieve low regret for small grids, DYS-net model achieves a low regret for all grids. In addition to training faster, DYS-net can also be trained for much larger problems, e.g., 100-by-100 grids, as shown in Figure 4. We found that CVX-net could not handle grids larger than 30-by-30, i.e. , problems with more than 1740 variables [^7] This is to be expected, as discussed in in Amos & Kolter (2017); Agrawal et al. (2019a) (see Table 1). Importantly, PertOpt-net takes close to a week to train for the 100-by-100 problem, whereas DYS-net takes about a day (see right Figure 4b). On the other hand, the training speed of BB-net is comparable to that of DYS-net, but does not lead to competitive accuracy as shown in Figure 4(a). Interpreting the outputs of DYS-net and CVX-net as (unnormalized) probabilities over the grid, one can use a greedy algorithm to determine the most probable path from top-left to bottom-right. For small grids, e.g. 5-by-5, this path coincides exactly with the true path for most $d$ (see Fig. 3).
5.3 Warcraft shortest path prediction
Finally, as an illustrative example, we consider the Warcraft terrains dataset first studied in Vlastelica et al. (2019). As shown in Figure 1, $d$ is a 96-by-96 RGB image, divided into a 12-by-12 grid of terrain tiles. Each tile has a different cost to traverse, and the goal is to find the quickest path from the top-left corner to the bottom-right corner.
Models, Training, and Evaluation
We build upon the code provided as part of the PyEPO package Tang & Khalil (2022). We use the same architecture for $w_{\Theta}(d)$ for BB-net, PertOpt-net, and DYS-net—a truncated ResNet18 He et al. (2016) as first proposed in Vlastelica et al. (2019). The baseline is the same truncated ResNet18 architecture trained to predict $x^{\star}(d)$ directly. We train each network for 50 epochs, except for the baseline, which is trained for 150 epochs. The initial learning rate is $5\times 10^{-4}$ and it is decreased by a factor of 10 after 30 and 40 epochs respectively. PertOpt-net and DYS-net are trained to minimize the $\ell_{2}$ loss (5), while BB-net is trained to minimize the Hamming loss as described in Vlastelica et al. (2019). Given test data $\{(d_{i},w(d_{i}),x^{\star}(d_{i}))\}_{i=1}^{N}$, we follow Tang & Khalil (2022) and evaluate a fully trained model $w_{\Theta}(d)$ using relative regret:
$\hat{r}=\sum_{i=1}^{N}\frac{r(d_{i})}{w(d_{i})^{\top}x^{\star}(d_{i})}$ (28)
Results
The results for this experiment are shown in Table 2. Interestingly, BB-net and PertOpt-net are much more competitive in this experiment than in the experiements of Sections 5.1.1, 5.1.2, and 5.2. We attribute this to the “discrete” nature of the true cost vector $w(d)$—for the Warcraft problem entries of $w(d)$ can only take on four, well-separated values—as opposed to the more “continuous” nature of $w(d)$ in the previous experiments. This difference is illustrated in Figure 5.
Figure 5: Left two figures: Sample cost matrices for shortest problem considered in Section 5.1.1. Right two figures: Sample cost matrices for the Warcraft shortest path prediction problem considered in Section 5.3. Note that in Section 5.3 the node weighted shortest path prediction problem is considered, while in Section 5.1.1 the edge weighted variant is solved. For ease of comparison, in the left two figures we have reshaped the edge cost vector into a node cost matrix.
Algorithm Test Relative Regret Prediction Accuracy Time (in hours)
Baseline $19.1150$ $0.0\%$ $0.07$
BB-net $0.1435$ $91.43\%$ $5.56$
PertOpt-net $0.4566$ $35.71\%$ $3.15$
DYS-net $0.2745$ $51.43\%$ $4.13$
Table 2: Results for the 12-by-12 Warcraft shortest path prediction problem. The low relative regret of BB-net could be attributed to the “discrete” nature of the true cost vectors as well as its use of the Hamming loss in training.
6 Conclusions
This work presents a new method for learning to solve ILPs using Davis-Yin splitting which we call DYS-net. We prove that the gradient approximation computed in the backward pass of DYS-net is indeed a descent direction, thus advancing the current understanding of implicit networks. Our experiments show that DYS-net is capable of scaling to truly large ILPs, and outperforms existing state-of-the-art methods in some settings. Our experiments also reveal an interesting dichotomy between problems in which entries of $w(d)$ may take on only a handful of discrete values and problems in which $w(d)$ is more “continuous”. Future work could explore this dichotomy further, as well as apply DYS-net to additional ILPs, for example the traveling salesman problem.
Appendix A Appendix A: Proofs
For the reader’s convenience we restate each result given in the main text before proving it.
Theorem 2. Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be as in (14), and suppose $f_{\Theta}(x;\gamma,d)=w_{\Theta}(d)^{\top}x+\frac{\gamma}{2}\|x\|_{2}^{2}$ for any neural network $w_{\Theta}(d)$. For any $\alpha\in(0,2/\gamma)$ define the sequence $\{z^{k}\}$ by:
$z^{k+1}=T_{\Theta}(z^{k})\quad\text{ for all }k\in\mathbb{N}$ (29)
where
$T_{\Theta}(z)\!\!\triangleq z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\!\left(\left(2-\alpha\gamma\right)P_{\mathcal{C}_{2}}(z)\!-\!z\!-\!\alpha w_{\Theta}(d)\right).$ (30)
If $x^{k}\triangleq P_{\mathcal{C}_{2}}(z^{k})$ then $\|x^{k+1}-x^{k}\|^{2}_{2}=O(1/k)$.
Proof.
First note that $\nabla_{x}f_{\Theta}({x}_{\Theta};\gamma,d)=w_{\Theta}(d)+\gamma x$. Because $f_{\Theta}({x}_{\Theta};\gamma,d)$ is strongly convex, ${x}_{\Theta}$ is unique and is characterized by the first order optimality condition:
$\nabla_{x}f_{\Theta}({x}_{\Theta};d)^{\top}\left(x-{x}_{\Theta}\right)\geq 0\quad\text{ for all }x\in\mathcal{C}.$ (31)
Note that (31) can equally be viewed as a variational inequality with operator $F_{\Theta}(x;d)=\nabla_{x}f_{\Theta}({x}_{\Theta};d)$. We deduce that $z_{\Theta}$ with $x_{\Theta}=P_{\mathcal{C}_{2}}(z_{\Theta})$ is a fixed point of
$T_{\Theta}(z)$ $\triangleq z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\left(2\cdot P_{\mathcal{C}_{2}}(z)\!-\!z\!-\!\alpha\left[w_{\Theta}(d)+\gamma P_{\mathcal{C}_{2}}(z)\right])\right)$
$=z\!-\!P_{\mathcal{C}_{2}}(z)+P_{\mathcal{C}_{1}}\left(\left(2-\alpha\gamma\right)\cdot P_{\mathcal{C}_{2}}(z)-z-\alpha w_{\Theta}(d)\right)$ (32a)
from (McKenzie et al., 2021, Theorem 4.2), which is itself a standard application of ideas from operator splitting Davis & Yin (2017); Ryu & Yin (2022).
As $\nabla_{x}f_{\Theta}({x}_{\Theta};\gamma,d)$ is $\gamma$-Lipschitz continuous it is also $\nabla_{x}f_{\Theta}$ is $1/\gamma$-cocoercive by the Baillon-Haddad theorem Baillon & Haddad (1977); Bauschke & Combettes (2009). (McKenzie et al., 2021, Theorem 3.3) then implies that $z^{k}\to z_{\Theta}$ but does not give a convergence rate. In fact, the convergence rate can be deduced from known results; because $T_{\Theta}$ is averaged for $\alpha<2/\gamma$ the rate follows from (Ryu & Yin, 2022, Theorem 1). ∎
Next, we state two auxiliary lemmas relating the Jacobian matrices to projections onto linear subspaces.
Lemma 8.
If $\mathcal{C}_{1}\triangleq\{x:Ax=b\}$, for full-rank $A\in{\mathbb{R}}^{m\times n}$, $b\in{\mathbb{R}}^{n}$ with $m<n$, and $\mathcal{H}_{1}\triangleq\operatorname{Null}(A)$ then
$\frac{\partial P_{\mathcal{C}_{1}}}{\partial z}=P_{\mathcal{H}_{1}},\quad\mbox{for all}\ z\in{\mathbb{R}}^{n}.$ (33)
Proof.
Let $A=U\Sigma V^{\top}$ denote the reduced SVD of $A$, and note that as $A\in\mathbb{R}^{m\times n}$ with $m<n$ we have $U\in\mathbb{R}^{m\times m},\Sigma\in\mathbb{R}^{m\times m}$ and $V\in\mathbb{R}^{n\times m}$. Differentiating the formula for $P_{\mathcal{C}_{1}}$ given in Lemma 1 yields
$\frac{\partial P_{\mathcal{C}_{1}}}{\partial z}=I-A^{\dagger}A,$ (34)
where $A^{\dagger}\triangleq V\Sigma^{-1}U^{\top}$. Note
$A^{\dagger}A=\left(V\Sigma^{-1}U^{\top}\right)\left(U\Sigma V^{\top}\right)=VV^{\top},$ (35)
which is the orthogonal projection onto $\operatorname{Range}(V)=\operatorname{Range}(A^{\top})$. It follows that $I-A^{\dagger}A$ is the orthogonal projection on to $\operatorname{Range}(A^{\top})^{\bot}=\operatorname{Null}(A)$. ∎
Lemma 9.
Define the multi-valued function
$c(\alpha)\triangleq\partial\max(0,\alpha)=\begin{cases}1&\text{ if }\alpha>0\\ 0&\text{ if }\alpha<0\\ [0,1]&\text{ if }\alpha=0\end{cases}$ (36)
and, for $z\in{\mathbb{R}}^{n}$, define $\mathcal{H}_{2,z}\triangleq\operatorname{Span}(e_{i}:z_{i}>0)$. Then
$\partial P_{\mathcal{C}_{2}}(\bar{z})=\left[\frac{\mathrm{d}}{\mathrm{d}z}\operatorname{ReLU}(z)\right]_{z=\overline{z}}=\mathrm{diag}(c(\overline{z})),$ (37)
and adopting the convention for choosing an element of $\partial P_{\mathcal{C}_{2}}(\bar{z})$ stated in the main text:
$\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z=\bar{z}}=\mathrm{diag}(\tilde{c}(\overline{z}))=P_{\mathcal{H}_{2,\overline{z}}}.$ (38)
Proof.
First, suppose $\overline{z}\in{\mathbb{R}}^{n}$ satisfies $\overline{z}_{i}\neq 0$, for all $i\in[n]$, i.e. $\overline{z}$ is a smooth point of $P_{\mathcal{C}_{2}}$. Note
$\frac{\mathrm{d}[\operatorname{ReLU}(\overline{z}_{i})]}{\mathrm{d}z}=1\ \mbox{if}\ i=j\ \mbox{and}\ z_{i}>0\qquad\mbox{and}\qquad\frac{\mathrm{d}[\operatorname{ReLU}(\overline{z}_{i})]}{\mathrm{d}z}=0\ \mbox{if}\ i\neq j\ \mbox{or}\ z_{i}<0.$ (39)
Thus, the Jacobian matrix is diagonal with a $1$ in the $(i,i)$-th position whenever $\overline{z}_{i}>0$ and 00 otherwise, i.e. $\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z=\bar{z}}=\mathrm{diag}(c(\overline{z}))$. Now suppose $z_{i}=0$ for one $i$. For all $z\in{\mathbb{R}}^{n}$ with $z_{i}<0$, the Jacobian $\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z}$ is well-defined and has a 0 in the $(i,i)$-th position, while for $z\in{\mathbb{R}}^{n}$ with $z_{i}>0$, the Jacobian $\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z}$ is well-defined and has a 1 in the $(i,i)$-th position. Taking the convex hull yields the interval $[0,1]$ in the $(i,i)$-th position, as claimed. The case where $\overline{z}_{i}=0$ for multiple $i$ is similar.
Consequently, the product of $\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z=\bar{z}}$ and any vector $v\in{\mathbb{R}}^{n}$ equals $v$ if and only if $v\in\operatorname{Span}(e_{i}:\overline{z}_{i}>0)$. This shows the linear operator is idempotent with fixed point set $\mathcal{H}_{2,\overline{z}}$, i.e. it is the projection operator $P_{\mathcal{H}_{2,\overline{z}}}$. ∎
Theorem 4. Suppose $A$ is full-rank and $\mathcal{H}_{1}\triangleq\operatorname{Null}(A)$, $\mathcal{H}_{2,z}\triangleq\operatorname{Span}\left(e_{i}:z_{i}>0\right)$. Then for all $\hat{z}\in{\mathbb{R}}^{n}$
$\left.\frac{\partial T_{\Theta}}{\partial z}\right|_{z=\bar{z}}=P_{\mathcal{H}_{1}^{\bot}}P_{\mathcal{H}_{2,\bar{z}}^{\bot}}+(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\bar{z}}}.$ (40)
Proof.
Differentiating the expression for $T_{\Theta}$ in (30) with respect to $z$ yields
$\left.\frac{\partial T_{\Theta}}{\partial z}\right|_{z=\hat{z}}$ $=I-\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z=\hat{z}}+\left.\frac{\mathrm{d}P_{\mathcal{C}_{1}}}{\mathrm{d}z}\right|_{z=y(\hat{z})}\left[(2-\alpha\gamma)\cdot\left.\frac{\mathrm{d}P_{\mathcal{C}_{2}}}{\mathrm{d}z}\right|_{z=\hat{z}}-I\right]$ (41a)
$=I-P_{\mathcal{H}_{2,\hat{z}}}+P_{\mathcal{H}_{1}}\left((2-\alpha\gamma)P_{\mathcal{H}_{2,\hat{z}}}-I\right),\quad\mbox{for all}\ \hat{z}\in{\mathbb{R}}^{n},$ (41b)
where, for notational brevity, we set $y(\hat{z})\triangleq\left(2-\alpha\gamma\right)\cdot P_{\mathcal{C}_{2}}(\hat{z})-\hat{z}-\alpha w_{\Theta}(d)$ in the first line and the second line follows from Lemmas 8 and 9. Repeatedly using the fact, for any subspace $\mathcal{H}\subset\mathbb{R}^{n}$, $P_{\mathcal{H}^{\bot}}=I-P_{\mathcal{H}}$, the derivative $\partial T_{\Theta}/\partial z$ can be further rewritten:
$\left.\frac{\partial T_{\Theta}}{\partial z}\right|_{z=\hat{z}}$ $={\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}\pgfsys@color@rgb@stroke{0}{0.5}{0}\pgfsys@color@rgb@fill{0}{0.5}{0}I-P_{\mathcal{H}_{2,\hat{z}}}}+(2-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}}-P_{\mathcal{H}_{1}}$ (42a)
$={\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}\pgfsys@color@rgb@stroke{0}{0.5}{0}\pgfsys@color@rgb@fill{0}{0.5}{0}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}+(2-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}\left(I-P_{\mathcal{H}_{2,\hat{z}}^{\bot}}\right)-P_{\mathcal{H}_{1}}$ (42b)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}+P_{\mathcal{H}_{1}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}}-P_{\mathcal{H}_{1}}$ (42c)
$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}(I-P_{\mathcal{H}_{1}})P_{\mathcal{H}_{2,\hat{z}}^{\top}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}(I-P_{\mathcal{H}_{2,\hat{z}}^{\bot}})}$ (42d)
$=P_{\mathcal{H}_{1}^{\bot}}P_{\mathcal{H}_{2,\hat{z}}^{\bot}}+(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,\hat{z}}},\quad\mbox{for all}\ \hat{z}\in{\mathbb{R}}^{n},$ (42e)
completing the proof. ∎
We use the following lemma to prove Theorem 6.
Lemma 10.
If the LICQ condition holds at ${x}_{\Theta}$, then $\mathcal{H}_{1}^{\bot}\cap\mathcal{H}_{2,{z}_{\Theta}}^{\bot}=\{0\}$.
Proof.
We first rewrite $\mathcal{H}_{1}^{\perp}$ and $\mathcal{H}_{2,{z}_{\Theta}}^{\perp}$. The subspace $\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$ is spanned by all non-positive coordinates of ${z}_{\Theta}$. By (17), $[{x}_{\Theta}]_{i}=\max\{0,[{z}_{\Theta}]_{i}\}$, and so $i\in\mathcal{A}({x}_{\Theta})$ if and only if $[{z}_{\Theta}]_{i}\leq 0$. It follows that
$\mathcal{H}_{2,{z}_{\Theta}}^{\bot}\triangleq\operatorname{Span}\{e_{i}:[{z}_{\Theta}]_{i}\leq 0\}=\operatorname{Span}\{e_{i}:i\in\mathcal{A}({x}_{\Theta})\}=\operatorname{Span}\{e_{i_{1}},\ldots,e_{i_{\ell}}\},$ (43)
where we enumerate $\mathcal{A}(x_{\Theta})$ via $\mathcal{A}(x_{\Theta})=\{i_{1},\ldots,i_{\ell}\}$. On the other hand, $\mathcal{H}_{1}^{\bot}=\operatorname{Range}(A^{\top})=\operatorname{Span}(a_{1},\ldots,a_{m})$ where $a_{i}^{\top}$ denotes the $i$-th row of $A$.
Let $v\in\mathcal{H}_{1}^{\bot}\cap\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$ be given. Since $v\in\mathcal{H}_{1}^{\perp}$, there are scalars $\alpha_{1},\ldots,\alpha_{\ell}$ such that $v=\alpha_{1}e_{i_{1}}+\cdots+\alpha_{\ell}e_{i_{\ell}}.$ Similarly, since $v\in\mathcal{H}_{2,{z}_{\Theta}}^{\perp}$, there are scalars $\beta_{1},\ldots,\beta_{m}$ such that $v=\beta_{1}a_{1}+\cdots+\beta_{m}a_{m}.$ Hence
$0=v-v=\big{(}\alpha_{1}e_{i_{1}}+\ldots+\alpha_{\ell}e_{i_{\ell}}\big{)}-\big{(}\beta_{1}a_{1}+\ldots+\beta_{m}a_{m}\big{)}.$ (44)
By the LICQ condition, $\{e_{i_{1}},\ldots,e_{i_{\ell}}\}\cup\{a_{1},\ldots,a_{m}\}$ is a linearly independent set of vectors; hence $\alpha_{1}=\ldots=\alpha_{\ell}=\beta_{1}=\ldots=\beta_{m}=0$ and, thus, $v=0$. Since $v$ was arbitrarily chosen in $\mathcal{H}_{1}^{\bot}\cap\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$, the result follows. ∎
Theorem 6. If the LICQ condition holds at ${x}_{\Theta}$, $A$ is full-rank and $\alpha\in(0,2/\gamma)$, then $\|\partial T_{\Theta}/\partial z\|_{z={z}_{\Theta}}<1$.
Proof.
By Lemma 10, the LICQ condition implies $\mathcal{H}_{1}^{\bot}\cap\mathcal{H}_{2,{z}_{\Theta}}^{\bot}=\{0\}$. This implies that either (i) the first principal angle $\tau$ between these two subspaces is nonzero, and so the cosine of this angle is less than unity, i.e.
$1>\cos(\tau)\triangleq\max_{u\in\mathcal{H}_{1}^{\bot}:\|u\|=1}\max_{v\in\mathcal{H}_{2,z}^{\bot}:\|v\|=1}\langle u,v\rangle,$ (45)
or (ii) (at least) one of $\mathcal{H}_{1}^{\bot},\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$ is the trivial vector space $\{0\}$. In either case, let $w\in\mathbb{R}^{n}$ be given. By Theorem 4, in case (ii)
$\left[\frac{\partial T_{\Theta}}{\partial z}w\right]_{z={z}_{\Theta}}=P_{\mathcal{H}_{1}^{\bot}}P_{\mathcal{H}_{2,{z}_{\Theta}}^{\bot}}w+(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,{z}_{\Theta}}}w=(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,{z}_{\Theta}}}w$ (46)
implying that
$\left\|\frac{\partial T_{\Theta}}{\partial z}w\right\|_{z={z}_{\Theta}}=(1-\alpha\gamma)\left\|P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,{z}_{\Theta}}}w\right\|\leq(1-\alpha\gamma)\|w\|,$ (47)
where the inequality follows as projection operators are firmly nonexpansive. In case (i), write $w=w_{1}+w_{2}$, where $w_{1}\in\mathcal{H}_{2,{z}_{\Theta}}$ and $w_{2}\in\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$. Appealing to Theorem 4 again
$\left[\frac{\partial T_{\Theta}}{\partial z}w\right]_{z={z}_{\Theta}}=P_{\mathcal{H}_{1}^{\bot}}P_{\mathcal{H}_{2,{z}_{\Theta}}^{\bot}}w+(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}P_{\mathcal{H}_{2,{z}_{\Theta}}}w=P_{\mathcal{H}_{1}^{\bot}}w_{2}+(1-\alpha\gamma)\cdot P_{\mathcal{H}_{1}}w_{1}.$ (48)
Pythagoras’ theorem may be applied to deduce, together with the fact $P_{\mathcal{H}_{1}^{\bot}}w_{2}$ and $P_{\mathcal{H}_{1}}w_{1}$ are orthogonal,
$\left\|\frac{\partial T_{\Theta}}{\partial z}w\right\|^{2}_{z={z}_{\Theta}}=\left\|P_{\mathcal{H}_{1}^{\bot}}w_{2}\right\|^{2}+(1-\alpha\gamma)^{2}\cdot\left\|P_{\mathcal{H}_{1}}w_{1}\right\|^{2}.$ (49)
Since $w_{2}\in\mathcal{H}_{2,{z}_{\Theta}}^{\bot}$, the angle condition (45) implies
$\left\|P_{\mathcal{H}_{1}^{\bot}}w_{2}\right\|^{2}=\langle P_{\mathcal{H}_{1}^{\bot}}w_{2},P_{\mathcal{H}_{1}^{\bot}}w_{2}\rangle=\langle w_{2},P_{\mathcal{H}_{1}^{\bot}}P_{\mathcal{H}_{1}^{\bot}}w_{2}\rangle=\langle w_{2},P_{\mathcal{H}_{1}^{\bot}}w_{2}\rangle\leq\cos(\tau)\cdot\|w_{2}\|^{2},$ (50)
where the third equality holds since orthogonal linear projections are symmetric and idempotent. Because projections are non-expansive and $P_{\mathcal{H}_{2,{z}_{\Theta}}}$ is linear,
$\left\|P_{\mathcal{H}_{1}}w_{1}\right\|^{2}=\left\|P_{\mathcal{H}_{1}}w_{1}-P_{\mathcal{H}_{1}}0\right\|^{2}\leq\|w_{1}-0\|^{2}=\|w_{1}\|^{2}.$ (51)
Combining (49), (50) and (51) reveals
${\left\|\frac{\partial T_{\Theta}}{\partial z}w\right\|^{2}_{z={z}_{\Theta}}}$ $\leq\cos(\tau)\cdot\|w_{2}\|^{2}+(1-\alpha\gamma)^{2}\|w_{1}\|^{2}$ (52a)
$\leq\max\{\cos(\tau),(1-\alpha\gamma)^{2}\}\cdot\left(\|w_{1}\|^{2}+\|w_{2}\|^{2}\right)$ (52b)
$=\max\{\cos(\tau),(1-\alpha\gamma)^{2}\}\cdot\|w\|^{2},$ (52c)
noting the final equality holds since $w_{1}$ and $w_{2}$ are orthogonal. Because (A) holds for arbitrarily chosen $w\in{\mathbb{R}}^{n}$,
${\left\|\frac{\partial T_{\Theta}}{\partial z}\right\|_{z={z}_{\Theta}}}\triangleq\sup\left\{{\left\|\frac{\partial T_{\Theta}}{\partial z}w\right\|_{z=z_{\Theta}}}:\|w\|=1\right\}\leq\sqrt{\max\{\cos(\tau),(1-\alpha\gamma)^{2}\}}<1,$ (53)
where the final inequality holds by (45) and the fact $\alpha\in(0,2/\gamma)$ implies $1-\alpha\gamma\in(-1,1)$, as desired. ∎
Corollary 7. If $T_{\Theta}$ is continuously differentiable with respect to $\Theta$ at $z_{\Theta}$, the assumptions in Theorem 6 hold and $(\partial T_{\Theta}/\partial\Theta)^{\top}(\partial T_{\Theta}/\partial\Theta)$ has condition number sufficiently small, then $p_{\Theta}(x_{\Theta})$ is a descent direction for $\ell$ with respect to $\Theta$.
Proof.
From the proof of Theorem 6 we see that $T_{\Theta}$ is contractive with constant $\Gamma=\sqrt{\max\{\cos(\tau),(1-\alpha\gamma)^{2}\}}$ and so the main theorem of Fung et al. (2022), guaranteeing $p_{\Theta}$ is a descent direction, as long as the condition number of $(\partial T_{\Theta}/\partial\Theta)^{\top}(\partial T_{\Theta}/\partial\Theta)$ is less than $1/\Gamma$. ∎
Remark 11.
Similar guarantees, albeit with less restrictive assumptions on $\partial T_{\Theta}/\partial\Theta$, can be deduced from the results of the recent work Bolte et al. (2024).
Appendix B Derivation for Canonical Form of Knapsack Prediction Problem
For completeness, we explain how to transform the $k$-knapsack prediction problem into the canonical form (8), and how to derive the standardized representation of the constraint polytope $\mathcal{C}$. Recall that the $k$-knapsack prediction problem, as originally stated, is
$x^{\star}=\operatorname*{arg\,max}_{x\in\mathcal{X}}w^{\top}x\text{ where }\mathcal{X}=\{x\in\{0,1\}^{\ell}:\ Sx\leq c\}\text{ and }S=\begin{bmatrix}s_{1}&\cdots&s_{\ell}\end{bmatrix}\in\mathbb{R}^{k\times\ell}$ (54)
We introduce slack variables $y_{1},\ldots,y_{k}$ so that the inequality constraint $Sx\leq c$ becomes
$-Sx+c\geq 0$ $\Longrightarrow-Sx+c=y\text{ and }y\geq 0$
$\Longrightarrow\begin{bmatrix}S&I_{k}\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}=c$
We relax the binary constraint $x_{i}\in\{0,1\}$ to $0\leq x_{i}\leq 1$. We add additional slack variables $z_{1},\ldots,z_{\ell}$ to account for the upper bound:
$1-x_{i}\geq 0\Longrightarrow 1-x_{i}=z_{i}\text{ and }z_{i}\geq 0\Longrightarrow\begin{bmatrix}I_{\ell\times\ell}&0_{\ell\times k}&I_{\ell\times\ell}\end{bmatrix}\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\mathbf{1}$ (59)
Putting this together, define
$A=\begin{bmatrix}-S&-I_{k\times k}&0_{k\times\ell}\\ I_{\ell\times\ell}&0_{\ell\times k}&I_{\ell\times\ell}\end{bmatrix}\in\mathbb{R}^{(k+\ell)\times(2\ell+k)}\text{ and }b=\begin{bmatrix}-c\\ \mathbf{1}_{\ell}\end{bmatrix}\in\mathbb{R}^{k+\ell}$ (60)
Finally, redefine $x=\begin{bmatrix}x&y&z\end{bmatrix}^{\top}$ and $w=\begin{bmatrix}-w&\mathbf{0}_{k}&\mathbf{0}_{\ell}\end{bmatrix}$ (where we’re using $-w$ to switch the argmax to an argmin) and obtain:
$x^{\star}=\operatorname*{arg\,min}_{x\in\mathrm{Conv}(\mathcal{X})}w^{\top}(d)x+\gamma\|x\|_{2}^{2}\text{ where }\mathrm{Conv}(\mathcal{X})=\{x:Ax=b\text{ and }x\geq 0\}$ (61)
Appendix C Experimental Details
C.1 Additional Data Details for Knapsack Problem
As mentioned, we use PyEPO Tang & Khalil (2022) to generate the training data. Specifically, $d\in{\mathbb{R}}^{5}$ is sampled from the multivariate Gaussian distribution with mean 00 and covariance $I$. Then, $B\in{\mathbb{R}}^{n\times 5}$ is sampled where each $B_{ij}=+1$ with probability $0.5$ and $-1$ with probability $0.5$. The associated cost vector $w(d)$ is computed as
$[w(d)]_{i}=\left[\frac{1}{3.5^{\mathrm{deg}}}\left(\frac{1}{\sqrt{5}}(Bd)_{i}+3\right)^{\mathrm{deg}}+1\right]\cdot\epsilon_{ij}$
where $\mathrm{deg}=4$ and $\epsilon_{ij}$ is sampled uniformly from the interval $[0.5,1.5]$.
C.2 Additional Training Details for Knapsack Problem
For all models we use an initial learning rate of $10^{-3}$ and a scheduler that reduces the learning rate whenever the validation loss plateaus. We also used weight decay with a parameter of $5\times 10^{-4}$. All networks were trained using a AMD Threadripper Pro 3955WX: 16 cores, 3.90 GHz, 64 MB cache, PCIe 4.0 CPU and an NVIDIA RTX A6000 GPU.
C.3 Additional Model Details for Large-Scale Shortest Path Problem
Our implementation of PertOpt-net used a PyTorch implementation [^8] See code at github.com/tuero/perturbations-differential-pytorch of the original TensorFlow code [^9] See code at github.com/google-research/google-research/tree/master/perturbations associated to the paper Berthet et al. (2020). We experimented with various hyperparameter settings for 5-by-5 grids and found setting the number of samples equal to 3, the temperature (i.e. $\varepsilon$) to 1 and using Gumbel noise to work best, so we used these values for all other shortest path experiments. Our implementation of BB-net uses the blackbox-backprop package [^10] See code at https://github.com/martius-lab/blackbox-backprop associated to the paper Pogančić et al. (2019). We found setting $\lambda=5$ (see Pogančić et al. (2019) for a description of this hyperparameter) to work best for 5-by-5 grids, so we use this value for all other shortest path experiments.
C.4 Additional Training Details for Shortest Path
We use the MSE loss to train DYS-net, PertOpt-net, and CVX-net. We tried using the MSE loss with BB-net but this did not work well, so we used the Hamming (also known as 0–1) loss, as done in Pogančić et al. (2019).
To train DYS-net and CVX-net, we use an initial learning rate of $10^{-2}$ and use a scheduler that reduces the learning rate whenever the test loss plateaus—we found this to perform the best for these two models. For PertOpt-net we found that using a fixed learning rate of $10^{-2}$ performed the best. For BB-net, we performed a logarithmic grid-search on the learning rate between $10{-1}$ to $10^{-4}$ and found that $10^{-3}$ performed best; we also attempted adaptive learning rate schemes such as reducing learning rates on plateau but did not obtain improved performance.
All networks were trained using a AMD Threadripper Pro 3955WX: 16 cores, 3.90 GHz, 64 MB cache, PCIe 4.0 CPU and an NVIDIA RTX A6000 GPU.
Appendix D Additional Experimental Results
In Figure 6, we show the test loss and training time per epoch for all three architectures: DYS-net, CVX-net, and PertOpt-net for 10-by-10, 20-by-20, and 30-by-30 grids. In terms of MSE loss, CVX-net and DYS-net lead to comparable performance. In the second row of Figure 6, we observe the benefits of combining the three-operator splitting with JFB Fung et al. (2022); in particular, DYS-net trains much faster. Figure 7 shows some randomly selected outputs for the three architectures once fully trained.
10-by-10 20-by-20 30-by-30
Test MSE Loss
Refer to caption Refer to caption Refer to caption
Training Time (in minutes)
Refer to caption Refer to caption Refer to caption
Figure 6: Comparison of of DYS-Net, cvxpylayers Agrawal et al. (2019a), PertOptNet Berthet et al. (2020), and Blackbox Backpropagation-net (BB-Net) Pogančić et al. (2019) for three different grid sizes: $10\times 10$ (first column), $20\times 20$ (second column), and $30\times 30$ (third column). The first row shows the MSE loss vs. epochs of the testing dataset. The second row shows the training time vs. epochs.
True Path DYS-net CVX-net PertOpt-net
10-by-10
Refer to caption Refer to caption Refer to caption Refer to caption
20-by-20
Refer to caption Refer to caption Refer to caption Refer to caption
30-by-30
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: True paths (column 1), paths predicted by DYS-net (column 2), CVX-net (column 3), and PertOpt-net (column 4). Samples are taken from different grid sizes: 10-by-10 (row 1), 20-by-20 (row 2), and 30-by-30 (row 3).